Bettina's age is three times Melvina's. If 20 is added to Melvina's age and 20 is subtracted from Bettina's their ages will be equal. How old is each now? Thank you so much for answering this question....God bless you..

Answer by pie90010(44)

3M = B
M + 20 = B - 20
This can be simplified to
3M - B = 0
-M + B = 40
+B and -B cancel each other out
2M = 40
M = 20
once you have M just plug it in to the original equation
3(20) = B
B = 60
Bettina is 60 and Melvina is 20